OpenStudy (anonymous):

If f(x) = x + 4, find the value of x if f(x) = 6.

OpenStudy (abb0t):

That means x+4=6 solve for x. do you know what to do now?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

we we know that f(x)=x +4 we also know that f(x)=6 so just equate the values of f(x)